OpenStudy (anonymous):

The circumference of a circle is 4x3.14 m^2n^4 units. What is the area of the circle? A.4m^4 n^16 square units B. 4x3.14^2 m^4 n^8 square units C. 4x3.14 m^4 n^8 square units D. 4x3.14 m^4 n^16 square units

OpenStudy (anonymous):

2 x 3.14 x r = 2 x 3.14 x 2m^2n^4 so r = 2m^2n^4 area = 3.14 * r^2 = 3.14 * (2m^2n^4 )^2 = 3.14 x 4m^4n^8 C. is correct